# Functional Regression with Nonstationarity and Error Contamination: Application to the Economic Impact of Climate Change

### v1

@@ C.1.1 Sample counterparts of $\theta_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ for feasible inference
Note that $\theta_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$, given in Theorems 4.3 and 4.4, are unknown, and thus the asymptotic results stated therein cannot be directly used for inference in practice. However, these unknown quantities can be replaced by reasonable estimators, which makes the results more useful in practice. To state the desired results, we introduce some additional notation. Let
@@
$(\widehat{D}_{\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}=\sum_{j=d_{N}+1}^{{\mathsf{K}}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{D}_{\kappa}]\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}]=\sum_{j=1}^{{\mathsf{K}}_{S}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{D}_{\kappa}^{S}]\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}^{S}].$
@@
In Corollary C.1, we use the assumptions employed for Theorem 4.3, including the assumption of distinct eigenvalues, (20). However, as discussed in Remark 4.4, this condition can be replaced by the two conditions given in Remark 4.4, allowing for the repetition of an eigenvalue; see the proof of Corollary C.1 in Section D.3.
@@ C.1.2 Pointwise asymptotic normality in the model with an intercept
In this section, we consider the model and estimator briefly discussed in Section 4.4 and extend the statistical inference methods developed in Section 4.3 to this case. We first introduce a set of assumptions, adapted from those in the previous sections. To this end, let $D_{c,\kappa}^{S}=(C^{S}_{c,\kappa})^{\ast}C^{S}_{c,\kappa}$ and $E_{c,\kappa}^{S}=C^{S}_{c,\kappa}(C^{S}_{c,\kappa})^{\ast}$, where $C^{S}_{c,\kappa}=\mathbb{E}[({\rm P}^{S}x_{t-\kappa}-\mu_{x,S})\otimes({\rm P}^{S}x_{t}-\mu_{x,S})]$ and $\mu_{x,S}=\mathbb{E}[{\rm P}_{S}x_{t}]$ ($\mathbb{E}[{\rm P}_{S}x_{t-\kappa}]$ due to stationarity). We also define $\theta_{c,{\mathsf{K}}_{S}}(\zeta)=\langle\zeta,(D_{\kappa}^{S})_{c,{\mathsf{K}}_{S}}^{-1}(C_{c,\kappa}^{S})^{\ast}\widetilde{C}_{c,0}^{S}C_{c,\kappa}^{S}(D_{c,\kappa}^{S})_{c,{\mathsf{K}}_{S}}^{-1}(\zeta)\rangle$, where $\widetilde{C}_{c,0}^{S}=\mathbb{E}[({\rm P}^{S}\tilde{x}_{t}-\mu_{x,S})\otimes({\rm P}^{S}\tilde{x}_{t}-\mu_{x,S})]$), and $\varpi_{c,t}(j,\ell)=\langle{\rm P}^{S}x_{t}-\mu_{x,S},\bm{\mathsf{v}}_{j}[D_{c,\kappa}^{S}]\rangle\langle{\rm P}^{S}x_{t-\kappa}-\mu_{x,S},\bm{\mathsf{v}}_{\ell}[E_{c,\kappa}^{S}]\rangle-\mathbb{E}[\langle{\rm P}^{S}x_{t}-\mu_{x,S},\bm{\mathsf{v}}_{j}[D_{c,\kappa}^{S}]\rangle\langle{\rm P}^{S}x_{t-\kappa}-\mu_{x,S},\bm{\mathsf{v}}_{\ell}[E_{c,\kappa}^{S}]\rangle]$.
@@
    The model (24) holds with Assumptions 1, 2 and E.
@@
$\widehat{\theta}_{c,{\mathsf{K}}_{S}}(\zeta)=\langle\zeta,(\widehat{D}_{c,\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}(\widehat{C}_{c,\kappa}^{S})^{\ast}\widehat{C}_{c,0}^{S}\widehat{C}_{c,\kappa}^{S}(\widehat{D}_{c,\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}(\zeta)\rangle\quad\text{and}\quad\widehat{C}_{c,t}=T^{-1}\sum_{t=1}^{T}\widehat{u}_{c,t}\otimes\widehat{u}_{c,t},$
where $\widehat{u}_{c,t}=y_{c,t}-\widehat{f}_{c,\kappa}(\tilde{x}_{c,t})$, $\widehat{C}_{c,\kappa}^{S}=\widehat{C}_{c,\kappa}\widehat{{\rm P}}^{S}_{c,\kappa}$, and $\widehat{C}_{c,0}^{S}=T^{-1}\sum_{t=1}^{T}\widehat{{\rm P}}^{S}_{c,\kappa}\tilde{x}_{c,t}\otimes\widehat{{\rm P}}^{S}_{\kappa}\tilde{x}_{c,t},$. The theoretical justification of this replacement is parallel to that in Section C.1.1, and will therefore be omitted.
@@ C.2 Robustness of the variance-ratio testing procedure for $d_{N}$
for some $d_{0}>0$. Let $\widehat{K}_{0}=T^{-1}\sum_{t=1}^{T}(\sum_{s=1}^{t}\tilde{x}_{t}\otimes\sum_{s=1}^{t}\tilde{x}_{t})$. We consider the variance-ratio (VR) test statistic, proposed by Nielsen et al., (2023) and further generalized by Nielsen et al., (2024), for examining (C.30). The test statistic is computed from the following eigenproblem:
@@
where $\widehat{{\rm P}}_{\ell}=\sum_{j=1}^{\ell}\bm{{\Pi}}_{j}[\widehat{C}_{0}]$ and $\ell\geq d_{0}$. The VR test statistic for examining (C.30) is then given by
@@
The robustness of the VR testing procedure to the existence of measurement errors is established by the following proposition:
@@
As discussed, for the consistency of the VR testing procedure, we need a conjectured upper bound $d_{\max}$ of $d_{N}$ and also $\ell\geq d_{N}$ (see Nielsen et al.,, 2023, Section 3.5). In the empirical study where this testing procedure is applied, we set $d_{\max}=\ell=5$.
@@ Appendix D Proofs
It will be convenient to introduce some notation in addition to that in Section 4.1. We define
$\widehat{E}_{\kappa}=\widehat{C}_{\kappa}\widehat{C}_{\kappa}^{\ast}.$
Similar to $\widehat{D}_{\kappa}$, $\widehat{E}_{\kappa}$ allows the following spectral decomposition:
@@
    As shown above, $T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S}$ is $o_{p}(1)$, and from this result, we note that $\|T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S}\|_{\textrm{op}}=\|(T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S})({\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S})\|_{\textrm{op}}=o_{p}(1)$. Thus we have
@@
Subsequently, we provide our proof of the desired result, focusing on the case where $\mathcal{H}_{y}=\mathcal{H}$, and hence $y_{t}$ and $u_{t}$ are function-valued, with $f$ understood as a bounded linear operator on $\mathcal{H}$. The other case, where $\mathcal{H}_{y}=\mathbb{R}$, is, as may be expected, simpler and requires only a trivial modification.
  • Proof of Theorem 4.2.
@@
    Define $\bm{\mathsf{v}}^{s}_{j}[D^{S}_{\kappa}]=\operatorname{sgn}\{\langle\bm{\mathsf{v}}_{j}[D^{S}_{\kappa}],\bm{\mathsf{v}}_{j}[\widehat{D}_{\kappa}^{S}]\rangle\}\bm{\mathsf{v}}_{j}[D^{S}_{\kappa}]$. We note that $\|\widehat{{\rm Q}}^{S}_{\kappa}\widehat{C}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-{C}_{\kappa}^{S}\|_{\textrm{op}}=\|\widehat{C}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-{C}_{\kappa}^{S}\|_{\textrm{op}}=O_{p}(T^{-1/2})$ and thus $\|\widehat{D}_{\kappa}^{S}-D^{S}_{\kappa}\|_{\textrm{op}}=\|\widehat{{\rm P}}^{S}_{\kappa}\widehat{D}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-D^{S}_{\kappa}\|_{\textrm{op}}=O_{p}(T^{-1/2})$ (see our proof of Theorem 4.3). Note also that
@@
    $f(\widehat{{\rm P}}_{\kappa}^{{\mathsf{K}}_{S}}-I){\rm P}^{S}(\zeta)=A_{1}+A_{2}+A_{3}+A_{4},$ (D.68)
@@ D.3 Proofs of the results in Section C
    Then the desired result i follows immediately. Moreover, by the consistency result $\|\widehat{f}_{\kappa}-f\|{\textrm{op}}\to_{p}0$ (Theorems 4.2 and 4.3), we obtain $\|\widehat{C}_{u}-\widetilde{C}_{u}\|_{\textrm{op}}\to_{p}0$, which establishes ii. $\square$
  • Proof of Corollary C.2.
@@ E.1 Spatial distribution of the regional growth rate
Having constructed the maximum possible coverage of regional income shares—and thereby gross regional product levels in a balanced panel (rectangular) form—we extend the series from 1951 to 2019 using two approaches, depending on sample length. Because regional income shares are relatively stable and tend to evolve gradually over time rather than exhibiting abrupt fluctuations, these smooth extrapolation methods are particularly appropriate. For regions with fewer than 15 available years, we extrapolate backward using the average ratio of the previous five years and forward using the average of the most recent five years (25 countries, 403 regions). For regions with 15 or more years of data, we fit the Nelson–Siegel two-factor model of Diebold and Li, (2006) to capture the gradual, nonlinear dynamics of income shares (56 countries, 1,238 regions).
@@
Figure 7 shows the time series of spatial distributions of regional growth rates, computed on the central 95% of the total probability mass using a nonparametric kernel estimator. [^7] Because the CLR transform requires strictly positive densities, we evaluate moments on the central 95% mass to avoid zero-density grids at the tails. The cross-sectional mean has trended downward over time. The cross-sectional variance rises until the early 1980s and then declines—consistent with the subsequent Great Moderation—albeit with temporary spikes around global downturns (e.g., the late 2000s). Since the mid-1980s, both skewness and kurtosis have increased, indicating a more right-skewed and leptokurtic distribution: while average growth has slowed, a subset of regions has experienced very high growth, thickening the right tail and increasing overall tail weight.

### v3

@@ C.1.1 Sample counterparts of $\theta_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$ for feasible inference
Note that $\theta_{{\mathsf{K}}_{S}}$ and $\widetilde{C}_{u}$, given in Theorems 4.3 and 4.4, are unknown, and thus the asymptotic results stated therein cannot be directly used for inference in practice. However, these unknown quantities can be replaced by reasonable estimators, which makes the results more useful in practice. To obtain the desired results, we introduce some additional notation. Let
@@
$(\widehat{D}_{\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}=\sum_{j=d_{N}+1}^{{\mathsf{K}}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{D}_{\kappa}]\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}]=\sum_{j=1}^{{\mathsf{K}}_{S}}\bm{\mathsf{\lambda}}^{-1}_{j}[\widehat{D}_{\kappa}^{S}]\bm{{\Pi}}_{j}[\widehat{D}_{\kappa}^{S}].$ (C.31)
@@
In Corollary C.1, we use the assumptions employed for Theorem 4.3, including the assumption of distinct eigenvalues, (19). However, as discussed in Remark 4.4, this condition can be replaced by the two conditions given in Remark 4.4, allowing for the repetition of an eigenvalue; see the proof of Corollary C.1 given in Section D.3.
@@ C.1.2 Pointwise asymptotic normality in the model with an intercept
In this section, we consider the model and estimator briefly discussed in Section 4.4 and extend the inferential methods developed in Section 4.3 to this case. We first introduce a set of assumptions adapted from those in the previous sections. To this end, let $D_{c,\kappa}^{S}=(C^{S}_{c,\kappa})^{\ast}C^{S}_{c,\kappa}$ and $E_{c,\kappa}^{S}=C^{S}_{c,\kappa}(C^{S}_{c,\kappa})^{\ast}$, where $C^{S}_{c,\kappa}=\mathbb{E}[({\rm P}^{S}x_{t-\kappa}-\mu_{x,S})\otimes({\rm P}^{S}x_{t}-\mu_{x,S})]$ and $\mu_{x,S}={\rm P}^{S}\mu_{x}=\mathbb{E}[{\rm P}^{S}x_{t}]=\mathbb{E}[u_{t}^{S}]$, which is nonzero in general, while, in the considered model, $\mathbb{E}[u_{t}^{N}]=\mathbb{E}[{\rm P}^{N}\Delta x_{t}]=0$. We also define $\theta_{c,{\mathsf{K}}_{S}}(\zeta)=\langle\zeta,(D_{\kappa}^{S})_{c,{\mathsf{K}}_{S}}^{-1}(C_{c,\kappa}^{S})^{\ast}\widetilde{C}_{c,0}^{S}C_{c,\kappa}^{S}(D_{c,\kappa}^{S})_{c,{\mathsf{K}}_{S}}^{-1}(\zeta)\rangle$, where $\widetilde{C}_{c,0}^{S}=\mathbb{E}[({\rm P}^{S}\tilde{x}_{t}-\mu_{x,S})\otimes({\rm P}^{S}\tilde{x}_{t}-\mu_{x,S})]$), and $\varpi_{c,t}(j,\ell)=\langle{\rm P}^{S}x_{t}-\mu_{x,S},\bm{\mathsf{v}}_{j}[D_{c,\kappa}^{S}]\rangle\langle{\rm P}^{S}x_{t-\kappa}-\mu_{x,S},\bm{\mathsf{v}}_{\ell}[E_{c,\kappa}^{S}]\rangle-\mathbb{E}[\langle{\rm P}^{S}x_{t}-\mu_{x,S},\bm{\mathsf{v}}_{j}[D_{c,\kappa}^{S}]\rangle\langle{\rm P}^{S}x_{t-\kappa}-\mu_{x,S},\bm{\mathsf{v}}_{\ell}[E_{c,\kappa}^{S}]\rangle]$.
@@
    The model (24) in Section 4.4 holds with Assumptions 1, 2, and E, with $x_{t}$ and $u_{t}^{S}$ being replaced by $x_{t}-\mu_{x}$ and $u_{t}^{S}-\mu_{x,S}$, respectively.
@@
$\widehat{\theta}_{c,{\mathsf{K}}_{S}}(\zeta)=\langle\zeta,(\widehat{D}_{c,\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}(\widehat{C}_{c,\kappa}^{S})^{\ast}\widehat{C}_{c,0}^{S}\widehat{C}_{c,\kappa}^{S}(\widehat{D}_{c,\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}(\zeta)\rangle\quad\text{and}\quad\widehat{C}_{c,u}=T^{-1}\sum_{t=1}^{T}\widehat{u}_{c,t}\otimes\widehat{u}_{c,t},$
where $\widehat{u}_{c,t}=y_{c,t}-\widehat{f}_{c,\kappa}(\tilde{x}_{c,t})$, $\widehat{C}_{c,\kappa}^{S}=\widehat{C}_{c,\kappa}\widehat{{\rm P}}^{S}_{c,\kappa}$, $\widehat{C}_{c,0}^{S}=T^{-1}\sum_{t=1}^{T}\widehat{{\rm P}}^{S}_{c,\kappa}\tilde{x}_{c,t}\otimes\widehat{{\rm P}}^{S}_{\kappa}\tilde{x}_{c,t},$ and $(\widehat{D}_{c,\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}$ is defined as $(\widehat{D}_{\kappa}^{S})_{{\mathsf{K}}_{S}}^{-1}$ in (C.31) but with $\widehat{D}_{\kappa}$ being replaced by $\widehat{D}_{c,\kappa}$ (or $\widehat{D}_{\kappa}^{S}$ being replaced by $\widehat{D}_{c,\kappa}^{S}$). The theoretical justification of this replacement is parallel to that in Section C.1.1, and will therefore be omitted.
@@ C.2 Robustness of the variance-ratio testing procedure for $d_{N}$
for some $d_{0}>0$. Hereafter, let
$\widehat{K}_{0}=T^{-1}\sum_{t=1}^{T}(\sum_{s=1}^{t}\tilde{x}_{t}\otimes\sum_{s=1}^{t}\tilde{x}_{t}).$
We consider the variance-ratio (VR) test statistic, proposed by Nielsen et al., (2023) and further generalized by Nielsen et al., (2024), for examining (C.32). The test statistic is computed from the following eigenproblem:
@@
where $\{\gamma\}_{j\geq 1}$ are the eigenvalues ordered from smallest to largest, $\{\phi_{j}\}_{j\geq 1}$ are the corresponding eigenvectors, $\widehat{{\rm P}}_{\ell}=\sum_{j=1}^{\ell}\bm{{\Pi}}_{j}[\widehat{C}_{0}]$ and $\ell\geq d_{0}$. The VR test statistic for examining (C.32) is then given by
@@
The robustness of the VR testing procedure to the presence of measurement errors is established by the following proposition:
@@
As discussed, for the consistency of the VR testing procedure, we need a conjectured upper bound $d_{\max}$ of $d_{N}$ and also $\ell\geq d_{N}$ (see Nielsen et al.,, 2023, Section 3.5). In the empirical study in Section 5, where this testing procedure is applied, we set $d_{\max}=\ell=5$.
Remark C.2
In our proposed model, the stochastic trends of $y_{t}$ are explained by those of $x_{t}$ via $f^{N}$. As an extension, the VR testing procedure may be used as a diagnostic check to see if this holds in the given dataset. For example, if this is true, then we have $\widehat{u}_{t}^{N}\coloneqq y_{t}-\widehat{f}_{\kappa}^{N}(\tilde{x}_{t})=u_{t}^{N}+O_{p}(T^{-1})$ (where ${u}_{t}^{N}=y_{t}-{f}^{N}_{\kappa}(\tilde{x}_{t})=f^{S}(\tilde{x}_{t})+u_{t}$) due to Theorem 4.2, and hence the nonstationarity dimension associated with $u_{t}^{N}$ must equal zero. By a straightforward modification of Lemma C.1 and Proposition C.1, it is immediately deduced that, if the stochastic trends of $y_{t}$ are explained by those of $x_{t}$, then the VR testing procedure applied to $\widehat{u}_{t}^{N}$ needs to conclude that its nonstationarity dimension equals zero. If the model with an intercept discussed in Section 4.4 is considered as in our empirical analysis in Section 5.2.1, this diagnostic check is applied to $\widehat{u}_{t}^{N}\coloneqq{y}_{c,t}-\widehat{f}_{\kappa}^{N}(\tilde{x}_{c,t})$ using the VR testing procedure, adjusted for an intercept, discussed in Remark C.1. We applied this diagnostic check with $d_{\max}=\ell=3$. These values are smaller than those employed in Table 1 but this is natural since the number of stochastic trends is expected to be smaller in this case; when the same VR testing procedure applied to $\tilde{x}_{t}$ is applied to $y_{t}$, we obtain $p$-values for $d_{0}=5,\ldots,1$ of 0.3%, 1.3%, 6.3%, 62.6%, and 75.4%, strongly supporting the presence of two (or possibly three) stochastic trends, with the number of stochastic trends in $u_{t}^{N}$ not exceeding this. We obtained test statistics of $3458.5,820.95$ and $101.09$ for $d_{0}=3,2,$ and $1$, respectively, and the corresponding $p$-values are less than $0.1\%$ for $d_{0}=3$ and $2$, and $4.7\%$ for $d_{0}=1$, which strongly supports that $u_{t}^{N}$ does not have a stochastic trend.
@@ Appendix D Proofs
It will be convenient to introduce some notation in addition to that in Section 4.1. Note first that, similar to $\widehat{D}_{\kappa}$, $\widehat{E}_{\kappa}$ (defined in Assumption 4) allows the following spectral decomposition:
@@
As shown above, $T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S}$ is $o_{p}(1)$, and from this result, we note that $\|T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S}\|_{\textrm{op}}=\|(T^{-1}{\rm P}^{N}\widehat{C}_{\kappa}^{\ast}{\rm P}^{S})({\rm P}^{S}\widehat{C}_{\kappa}{\rm P}^{S})\|_{\textrm{op}}=o_{p}(1)$. Thus, we have
@@
Subsequently, we provide our proof of the desired result, focusing on the case where $\mathcal{H}_{y}=\mathcal{H}$, and hence $y_{t}$ and $u_{t}$ are function-valued, with $f$ understood as a bounded linear operator on $\mathcal{H}$. The other case, where $\mathcal{H}_{y}=\mathbb{R}$, is, as may be expected, simpler and requires only a moderate modification.
@@
Let $\bm{\mathsf{v}}^{s}_{j}[D^{S}_{\kappa}]=\operatorname{sgn}\{\langle\bm{\mathsf{v}}_{j}[D^{S}_{\kappa}],\bm{\mathsf{v}}_{j}[\widehat{D}_{\kappa}^{S}]\rangle\}\bm{\mathsf{v}}_{j}[D^{S}_{\kappa}]$. We note that $\|\widehat{{\rm Q}}^{S}_{\kappa}\widehat{C}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-{C}_{\kappa}^{S}\|_{\textrm{op}}=\|\widehat{C}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-{C}_{\kappa}^{S}\|_{\textrm{op}}=O_{p}(T^{-1/2})$ and thus $\|\widehat{D}_{\kappa}^{S}-D^{S}_{\kappa}\|_{\textrm{op}}=\|\widehat{{\rm P}}^{S}_{\kappa}\widehat{D}_{\kappa}\widehat{{\rm P}}^{S}_{\kappa}-D^{S}_{\kappa}\|_{\textrm{op}}=O_{p}(T^{-1/2})$ (see our proof of Theorem 4.3). Note also that
@@
$f(\widehat{{\rm P}}_{\kappa}^{{\mathsf{K}}_{S}}-I){\rm P}^{S}(\zeta)=A_{1}+A_{2}+A_{3}+A_{4},$
@@ D.3 Proofs of the results in Section C
Then the desired result i follows immediately. Moreover, by the consistency result $\|\widehat{f}_{\kappa}-f\|_{\textrm{op}}\to_{p}0$ (Theorems 4.2 and 4.3), we obtain $\|\widehat{C}_{u}-\widetilde{C}_{u}\|_{\textrm{op}}\to_{p}0$, which establishes ii. $\square$
@@ E.1 Spatial distribution of the regional growth rate
Having constructed the maximum possible coverage of regional income shares—and thereby GRP levels in a balanced panel (rectangular) form—we extend the series from 1951 to 2019 using two approaches, depending on sample length. Because regional income shares are relatively stable and tend to evolve gradually over time rather than exhibiting abrupt fluctuations, these smooth extrapolation methods are particularly appropriate. For regions with fewer than 15 available years, we extrapolate backward using the average ratio of the previous five years and forward using the average of the most recent five years (25 countries, 403 regions). For regions with 15 or more years of data, we fit the Nelson–Siegel two-factor model of Diebold and Li, (2006) to capture the gradual, nonlinear dynamics of income shares (56 countries, 1,238 regions).
@@
Figure 7 shows the time series of spatial distributions of regional growth rates, computed over the central 95% of the total probability mass using a nonparametric kernel estimation method. [^7] Because the CLR transform requires strictly positive densities, we evaluate moments on the central 95% mass to avoid zero-density grids at the tails. The cross-sectional mean has trended downward over time. The cross-sectional variance rises until the early 1980s and then declines—consistent with the subsequent Great Moderation—albeit with temporary spikes around global downturns (e.g., the late 2000s). Since the mid-1980s, both skewness and kurtosis have increased, indicating a more right-skewed and leptokurtic distribution: while average growth has slowed, a subset of regions has experienced very high growth, thickening the right tail and increasing overall tail weight.